# Geometric Transformations on Null Curves in the Anti-de Sitter $3$-Space

### v1

@@
Abstract.
We provide a geometric transformation on null curves in the anti-de Sitter $3$-space (${\rm AdS}$) which induces the Bäcklund transformation for the KdV equation. In addition, we show that this geometric transformation satisfies a suitable permutability theorem. We also illustrate how to implement it when the original null curve has constant bending.
Keywords: Anti-de Sitter Space, Geometric Transformations, KdV Equation, Null Curves.
Mathematics Subject Classification 2020: 37K35, 53B30.
Authors partially supported by PRIN projects PRIN 2017 “Real and Complex Manifolds: Topology, Geometry and Holomorphic Dynamics” (protocollo 2017JZ2SW5-004) and PRIN 2022 “Real and Complex Manifolds: Geometry and Holomorphic Dynamics” (protocollo 2022AP8HZ9) and by the GNSAGA of INDAM. The authors gratefully acknowledge the warm hospitality of the Department of Mathematics at the Politecnico di Torino and of the Texas Tech University Center in Sevilla.
1. Introduction
In a previous paper ([8]), we showed that, in the context of Lorentzian geometry on the anti-de Sitter $3$-space (${\rm AdS}$, for short) there are flows (referred to as LIEN flows) on null curves that induce bending evolution by any partial differential equation (PDE) in the KdV hierarchy. In particular, the lower order LIEN flow induces evolution by the Korteweg-De Vries (KdV) equation in the form
@@ 1. Introduction
Although similar flows exist on other Lorentzian space forms (see, for instance, [1, 6]), the specific case of ${\rm AdS}$ is special in the sense that the theory of null curves in ${\rm AdS}$ is closely related to the theory of star-shaped curves in the centro-affine plane. Indeed, null curves in ${\rm AdS}$ can be analyzed by combining together pairs of star-shaped curves in the centro-affine plane ([8]).
The KdV equation is a classical PDE that originated as a model to understand the propagation of waves on shallow water surfaces ([5]) and is a prototype of completely integrable evolution equations. As a completely integrable PDE the KdV equation has a rich structure including, not only infinitely many conservation laws, but also a Bäcklund transformation ([11]) which generates new solutions from old. In this paper, we use Tabachnikov’s transformation on star-shaped curves ([10]) and a similar transformation of Terng-Wu ([12]) to define a geometric transformation (referred to as the $\mathcal{T}$-transform) on null curves in ${\rm AdS}$ that corresponds to the Bäcklund transformation for their bendings.
We begin, in Section 2, by recalling the basic properties of null curves $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ without inflection points in the ${\rm SL}(2,\mathbb{R})$ model for ${\rm AdS}$. Such curves possess canonical parameterizations and a third order differential invariant $\kappa$, the bending (often known as curvature or torsion). In addition, they also have a spinor frame field $(F_{+},F_{-}):J\subseteq\mathbb{R}\longrightarrow{\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$ defined along them. The components $F_{\pm}$ of the spinor frame field along $\gamma$ are, precisely, the canonical central affine frame fields of two star-shaped curves $\eta_{\pm}$ in the centro-affine plane ˙$\dot{\mathbb{R}}^{2}=\mathbb{R}^{2}\setminus\{(0,0)\}$ with central affine curvatures ${\rm k}_{\pm}=\kappa\pm 1$, respectively ([8]).
@@
Next, in Section 4, after briefly recalling some basic facts about the Bäcklund transformation for the KdV equation ([11, 13]), we use the $\mathcal{T}$-transform to find, in Theorem 4.4, a geometric realization of it. In addition, such a transformation satisfies a permutability theorem as well. This geometric realization consists on applying the $\mathcal{T}$-transform to a solution of the LIEN flow ([8]).
@@
The results of this paper naturally suggest further directions for research. In particular, it is likely that these results (as well as those of [8]) can be reformulated in the context of holomorphic null curves in ${\rm PSL}(2,\mathbb{C})$. Since the seminal work of R. Bryant ([3]), it is known that the geometry of holomorphic null curves ([4, 7]) in ${\rm PSL}(2,\mathbb{C})$ is related via the Bryant’s correspondence to the geometry of immersed surfaces with constant mean curvature one (CMC $1$, for short) in the hyperbolic space $\mathbb{H}^{3}$. In this case, it is expected that the $\mathcal{T}$-transform corresponds to the Darboux transform for CMC $1$ surfaces in $\mathbb{H}^{3}$. It is also reasonable to ask if the stationary solutions of the holomorphic LIEN flow or its finite-gap solutions can be used to find new embedded CMC $1$ surfaces with a finite number of ends, in the spirit of [2].
2. Preliminaries
In this section we will collect the basic information about the ${\rm SL}(2,\mathbb{R})$ model for the anti-de Sitter $3$-space (${\rm AdS}$) and about the geometric features of null curves in ${\rm AdS}$. For a more detailed analysis we refer the reader to Sections 2 and 3 of [8], respectively.
2.1. The Anti-De Sitter $3$-Space
@@
$U\wedge V=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\wedge\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\in\Lambda^{2}(\mathbb{R}^{2,2})\,,$
and we say that a bivector $W\wedge Z$ of type $(-,0)$ is positive if
${\rm det}\begin{pmatrix}\langle U,W\rangle&\langle U,Z\rangle\\ \langle V,W\rangle&\langle V,Z\rangle\end{pmatrix}>0\,.$
The set of all positive bivectors of type $(-,0)$ is denoted by $\mathcal{N}_{+}$. This choice defines the notion of time orientation in $\mathbb{R}^{2,2}$.
@@
Consider the normal vector field $Q:X\in{\rm AdS}\longmapsto-2X\in\mathbb{R}^{2,2}$ and the orientation in ${\rm AdS}$ determined by the volume form $i_{Q}\Omega$. Given a null tangent vector $T\in T_{X}{\rm AdS}$, the bivector $X\wedge T$ is of type $(-,0)$. We define a time-orientation on ${\rm AdS}$ by declaring $T$ to be future-directed if $X\wedge T\in\mathcal{N}_{+}$.
2.2. Geometry of Null Curves
Let $J\subseteq\mathbb{R}$ be an open interval. A smooth immersed curve $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ is null if its velocity vector $\gamma^{\prime}(s)$ is a null (or, light-like) vector for each $s\in J$. In other words, if $\langle\gamma^{\prime}(s),\gamma^{\prime}(s)\rangle=0$ for all $s\in J$. A null curve is future-directed if the bivector $\gamma\wedge\gamma^{\prime}$ of type $(-,0)$ is positive, that is, if $\gamma\wedge\gamma^{\prime}\in\mathcal{N}_{+}$.
Let $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a future-directed null curve without inflection points (ie. such that $\gamma^{\prime}\wedge\gamma^{\prime\prime}\neq 0$ holds). Using the terminology of [8], we say that $\gamma$ is parameterized by the proper time if $\langle\gamma^{\prime\prime}(s),\gamma^{\prime\prime}(s)\rangle=4$ for every $s\in J$. The bending of a curve $\gamma$ parameterized by the proper time is defined by
$\kappa(s)=-\frac{1}{16}\langle\gamma^{\prime\prime\prime}(s),\gamma^{\prime\prime\prime}(s)\rangle\,.$
@@ 2.2. Geometry of Null Curves
From now on, we assume that all our curves $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ are null, future-directed, parameterized by their proper time $s\in J$, and have no inflection points.
Let $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$. The spinor frame field along $\gamma$ is a lift $(F_{+},F_{-}):J\subseteq\mathbb{R}\longrightarrow{\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$ such that
(2.3) $dF_{+}$ $=$ $F_{+}\begin{pmatrix}0&\kappa+1\\ 1&0\end{pmatrix}ds\,,$
(2.6) $dF_{-}$ $=$ $F_{-}\begin{pmatrix}0&\kappa-1\\ 1&0\end{pmatrix}ds\,.$
The spinor frame field is defined up to a sign. The equations (2.3)-(2.6) are the spinorial counterpart of the classical Frenet-type equations and, hence, we will refer to them as the spinorial Frenet-type equations of $\gamma$. Observe that the null curve $\gamma$ can be recovered from the spinor frame field $(F_{+},F_{-})$ as
$\gamma=F_{+}F_{-}^{-1}\,.$
@@
Theorem 2.3 (Theorem 3.3 of [8]).
Let $\gamma:J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and spinor frame field $(F_{+},F_{-})$ along it. Then, the first column vectors of $F_{\pm}$ form a pair of star-shaped curves $(\eta_{+},\eta_{-})$ in ˙$\dot{\mathbb{R}}^{2}$ with central affine curvatures [^1] Our notion of central affine curvature coincides with that of Terng-Wu ([12]) and it has the opposite sign of that of Pinkall ([9]).
${\rm k}_{+}=\kappa+1\,,\quad\quad\quad{\rm k}_{-}=\kappa-1\,,$
respectively.
Conversely, let $(\eta_{+},\eta_{-})$ be a pair of star-shaped curves with central affine curvatures ${\rm k}_{+}$ and ${\rm k}_{-}$ related by $({\rm k}_{+}-{\rm k}_{-})/2=1$, and canonical central affine frame fields $F_{+}=(\eta_{+},\eta_{+}^{\prime})$ and $F_{-}=(\eta_{-},\eta_{-}^{\prime})$, respectively. Then,
$\gamma=F_{+}F_{-}^{-1}\,,$
is a null curve in ${\rm AdS}$ with bending $\kappa=({\rm k}_{+}+{\rm k}_{-})/2$ and spinor frame field $(F_{+},F_{-})$ along it.
@@
Following the terminology introduced in [8], the pair of star-shaped curves $(\eta_{+},\eta_{-})$ will be referred as to the pair of cousins associated with the null curve $\gamma$.
We briefly recall here the case of closed null curves with constant bending (for more details, see Example 3.5 of [8]).
@@
Consider a closed null curve $\gamma:J=\mathbb{R}\longrightarrow{\rm AdS}$ with constant bending $\kappa$. Since $\gamma$ is closed, this implies that the spinor frame field $(F_{+},F_{-})$ along $\gamma$ is periodic and that the least periods of $F_{+}$ and $F_{-}$ are commensurable. That is, the bending of $\gamma$ must be
@@
where $m>n$ are two relatively prime natural numbers. Furthermore:
  1. (1)
@@
In Figure 1 we illustrate a closed null curve with constant bending as well as its associated pair of star-shaped cousins (other illustrations of closed null curves of this family can be found in Figures 2 and 3 of [8]).
Refer to caption   Refer to caption
Figure 1. Left: The closed null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Example 2.5). This curve represents a torus knot of type $(2,5)$. Right: The associated pair of star-shaped cousins $(\eta_{+},\eta_{-})$ in blue and red, respectively.
3. The $\mathcal{T}$-Transform
In this section we will introduce a geometric transformation on null curves in ${\rm AdS}$, which we call the $\mathcal{T}$-transform, and show how to build it using a solution of the Riccati equation.
In order to define the $\mathcal{T}$-transform on null curves we will employ Tabachnikov’s transformation [^2] In [10], a symmetric relation is introduced for circle maps by considering the continuous version of a relation for ideal polygons in the hyperbolic plane. Extending this relation to the lifts of these circle maps to $\mathbb{R}^{2}$, a transformation is defined. This transformation is, essentially, our $\mathcal{T}$-transform on star-shaped curves. on star-shaped curves ([10]) and translate it to null curves through the associated pair of cousins.

### v2

@@
[…]
Geometric Transformations on Null Curves in the Anti-de Sitter 3-Space
[…]
Geometric Transformations on Null Curves
in the Anti-de Sitter 3-Space
[…]
Emilio MUSSO a and Álvaro PÁMPANO b
[…]
E. Musso and A. Pámpano
[…]
a) Dipartimento di Scienze Matematiche, Politecnico di Torino,
[…]
Received September 24, 2024, in final form February 05, 2025; Published online February 12, 2025
[…]
We provide a geometric transformation on null curves in the anti-de Sitter $3$-space (AdS) which induces the Bäcklund transformation for the KdV equation. In addition, we show that this geometric transformation satisfies a suitable permutability theorem. We also illustrate how to implement it when the original null curve has constant bending.
[…]
anti-de Sitter space; geometric transformations; KdV equation; null curves
[…]
37K35; 53B30
1 Introduction
In a previous paper [8], we showed that, in the context of Lorentzian geometry on the anti-de Sitter $3$-space (${\rm AdS}$, for short) there are flows (referred to as LIEN flows) on null curves that induce bending evolution by any partial differential equation (PDE) in the KdV hierarchy. In particular, the first order LIEN flow induces evolution by the Korteweg–de Vries (KdV) equation in the form
@@ 1 Introduction
Although similar flows exist on other Lorentzian space forms (see, for instance, [4, 6]), the specific case of ${\rm AdS}$ is special in the sense that the theory of null curves in ${\rm AdS}$ is closely related to the theory of star-shaped curves in the centro-affine plane. Indeed, null curves in ${\rm AdS}$ can be analyzed by combining together pairs of star-shaped curves in the centro-affine plane [8].
The KdV equation is a classical PDE that originated as a model to understand the propagation of waves on shallow water surfaces [5] and is a prototype of completely integrable evolution equations. As a completely integrable PDE the KdV equation has a rich structure including, not only infinitely many conservation laws, but also a Bäcklund transformation [11] which generates new solutions from old. In this paper, we use Tabachnikov’s transformation on star-shaped curves [10] and a similar transformation of Terng–Wu [12] to define a geometric transformation (referred to as the $\mathcal{T}$-transform) on null curves in ${\rm AdS}$ that corresponds to the Bäcklund transformation for their bendings.
We begin, in Section 2, by recalling the basic properties of null curves $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ without inflection points in the ${\rm SL}(2,\mathbb{R})$ model for ${\rm AdS}$. Such curves possess canonical parameterizations and a third order differential invariant $\kappa$, the bending (often known as curvature or torsion). In addition, they also have a spinor frame field $(F_{+},F_{-})\colon J\subseteq\mathbb{R}\longrightarrow{\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$ defined along them. The components $F_{\pm}$ of the spinor frame field along $\gamma$ are, precisely, the canonical central affine frame fields of two star-shaped curves $\eta_{\pm}$ in the centro-affine plane ˙$\dot{\mathbb{R}}^{2}=\mathbb{R}^{2}\setminus\{(0,0)\}$ with central affine curvatures ${\rm k}_{\pm}=\kappa\pm 1$, respectively [8].
@@
Next, in Section 4, after briefly recalling some basic facts about the Bäcklund transformation for the KdV equation [11, 13], we use the $\mathcal{T}$-transform to find, in Theorem 4.4, a geometric realization of it. In addition, such a transformation satisfies a permutability theorem as well. This geometric realization consists on applying the $\mathcal{T}$-transform to a solution of the LIEN flow [8].
@@
The results of this paper naturally suggest further directions for research. In particular, it is likely that these results (as well as those of [8]) can be reformulated in the context of holomorphic null curves in ${\rm PSL}(2,\mathbb{C})$. Since the seminal work of R. Bryant [2], it is known that the geometry of holomorphic null curves [3, 7] in ${\rm PSL}(2,\mathbb{C})$ is related via the Bryant’s correspondence to the geometry of immersed surfaces with constant mean curvature one (CMC $1$, for short) in the hyperbolic space $\mathbb{H}^{3}$. In this case, it is expected that the $\mathcal{T}$-transform corresponds to the Darboux transform for CMC $1$ surfaces in $\mathbb{H}^{3}$. It is also reasonable to ask if the stationary solutions of the holomorphic LIEN flow or its finite-gap solutions can be used to find new embedded CMC $1$ surfaces with a finite number of ends, in the spirit of [1].
2 Preliminaries
In this section, we will collect the basic information about the ${\rm SL}(2,\mathbb{R})$ model for the anti-de Sitter $3$-space (${\rm AdS}$) and about the geometric features of null curves in ${\rm AdS}$. For a more detailed analysis, we refer the reader to [8, Sections 2 and 3].
2.1 The anti-de Sitter 3-space
@@
$\langle\langle U\wedge V,W\wedge Z\rangle\rangle=\det\begin{pmatrix}\langle U,W\rangle&\langle U,Z\rangle\\ \langle V,W\rangle&\langle V,Z\rangle\end{pmatrix}.$
A bivector $U\wedge V\in\Lambda^{2}\bigl{(}\mathbb{R}^{2,2}\bigr{)}$ is of type (,)(-,-)( - , - ) if the restriction of the above inner product to ${\rm span}\{U\wedge V\}$ is negative definite, and of type $(-,0)$ if this restriction is nonzero semi-negative definite and degenerate. We fix the bivector of type (,)(-,-)( - , - )
$U\wedge V=\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\wedge\begin{pmatrix}0&1\\ -1&0\end{pmatrix}\in\Lambda^{2}\bigl{(}\mathbb{R}^{2,2}\bigr{)},$
and we say that a bivector $W\wedge Z$ of type $(-,0)$ is positive if $\langle\langle U\wedge V,W\wedge Z\rangle\rangle>0$. The set of all positive bivectors of type $(-,0)$ is denoted by $\mathcal{N}_{+}$. This choice defines the notion of time orientation in $\mathbb{R}^{2,2}$.
@@
Consider the normal vector field $Q\colon X\in{\rm AdS}\longmapsto-2X\in\mathbb{R}^{2,2}$ and the orientation in ${\rm AdS}$ determined by the volume form $i_{Q}\Omega$. Given a null tangent vector $T\in T_{X}{\rm AdS}$, the bivector ${X\wedge T}$ is of type $(-,0)$. We define a time-orientation on ${\rm AdS}$ by declaring $T$ to be future-directed if ${X\wedge T\in\mathcal{N}_{+}}$.
2.2 Geometry of null curves
Let $J\subseteq\mathbb{R}$ be an open interval. A smooth immersed curve $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ is null if its velocity vector $\gamma^{\prime}(s)$ is a null (or, light-like) vector for each $s\in J$. In other words, if $\langle\gamma^{\prime}(s),\gamma^{\prime}(s)\rangle=0$ for all $s\in J$. A null curve is future-directed if the bivector $\gamma\wedge\gamma^{\prime}$ of type $(-,0)$ is positive, that is, if $\gamma\wedge\gamma^{\prime}\in\mathcal{N}_{+}$.
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a future-directed null curve without inflection points (i.e., such that $\gamma^{\prime}\wedge\gamma^{\prime\prime}\neq 0$ holds). Since $\gamma^{\prime}\wedge\gamma^{\prime\prime}\neq 0$, then $\gamma^{\prime\prime}$ is a space-like vector. Using the terminology of [8], we say that $\gamma$ is parameterized by the proper time if $\langle\gamma^{\prime\prime}(s),\gamma^{\prime\prime}(s)\rangle=4$ for every $s\in J$. The bending of a curve $\gamma$ parameterized by the proper time is defined by
$\kappa(s)=-\frac{1}{16}\bigl{\langle}\gamma^{\prime\prime\prime}(s),\gamma^{\prime\prime\prime}(s)\bigr{\rangle}.$
@@ 2.2 Geometry of null curves
From now on, we assume that all our curves $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ are null, future-directed, parameterized by their proper time $s\in J$, and have no inflection points. In this setting, the bending $\kappa(s)$ completely determines the null curve $\gamma$, up to congruence.
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$. The spinor frame field along $\gamma$ is a lift $(F_{+},F_{-})\colon J\subseteq\mathbb{R}\longrightarrow{\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$ such that
${\rm d}F_{+}=F_{+}\begin{pmatrix}0&\kappa+1\\ 1&0\end{pmatrix}{\rm d}s,$ (2.1)
${\rm d}F_{-}=F_{-}\begin{pmatrix}0&\kappa-1\\ 1&0\end{pmatrix}{\rm d}s.$ (2.2)
The spinor frame field is defined up to a sign. The equations (2.1) and (2.2) are the spinorial counterpart of the classical Frenet-type equations and, hence, we will refer to them as the spinorial Frenet-type equations of $\gamma$. Observe that the null curve $\gamma$ can be recovered from the spinor frame field $(F_{+},F_{-})$ as $\gamma=F_{+}F_{-}^{-1}$.
Remark 2.3.
The frame $(F_{+},F_{-})$ can be explicitly obtained as a lifting to ${\rm SL}(2,\mathbb{R})\times{\rm SL}(2,\mathbb{R})$ of a moving frame adapted to the curve $\gamma$, with values in the automorphism group of ${\rm AdS}$. For the details of the construction, we refer to [8].
@@
Theorem 2.4 ([8, Theorem 3.3]).
Let $\gamma\colon J\subseteq\mathbb{R}\longrightarrow{\rm AdS}$ be a null curve with bending $\kappa$ and spinor frame field $(F_{+},F_{-})$ along it ((((which can be obtained by solving the spinorial Frenet-type equations (2.1) and (2.2))))). Then, the first column vectors of $F_{\pm}$ form a pair of star-shaped curves ${(\eta_{+},\eta_{-})}$ in ˙$\dot{\mathbb{R}}^{2}$ with central affine curvatures [^1] Our notion of central affine curvature coincides with that of Terng–Wu [12] and it has the opposite sign of that of Pinkall [9]. ${\rm k}_{+}=\kappa+1$, ${\rm k}_{-}=\kappa-1$, respectively.
Conversely, let $(\eta_{+},\eta_{-})$ be a pair of star-shaped curves with central affine curvatures ${\rm k}_{+}$ and ${\rm k}_{-}$ related by $({\rm k}_{+}-{\rm k}_{-})/2=1$, and canonical central affine frame fields $F_{+}=\bigl{(}\eta_{+},\eta_{+}^{\prime}\bigr{)}$ and ${F_{-}=\bigl{(}\eta_{-},\eta_{-}^{\prime}\bigr{)}}$, respectively. Then, $\gamma=F_{+}F_{-}^{-1}$, is a null curve in ${\rm AdS}$ with bending $\kappa=({\rm k}_{+}+{\rm k}_{-})/2$ and spinor frame field $(F_{+},F_{-})$ along it.
@@
Following the terminology introduced in [8], the pair of star-shaped curves ${(\eta_{+},\eta_{-})}$ will be referred to as the pair of cousins associated with the null curve $\gamma$.
We briefly recall here the case of closed null curves with constant bending (for more details, see [8, Example 3.5]).
@@
Consider a closed null curve $\gamma\colon J=\mathbb{R}\longrightarrow{\rm AdS}$ with constant bending $\kappa$. Since $\gamma$ is closed, this implies that the spinor frame field $(F_{+},F_{-})$ along $\gamma$ is periodic and that the least periods of $F_{+}$ and $F_{-}$ are commensurable. That is, the bending of $\gamma$ must be
@@
where $m>n$ are two relatively prime natural numbers. Furthermore,
  1. 1.
@@
In Figure 1 we illustrate a closed null curve with constant bending as well as its associated pair of star-shaped cousins (other illustrations of closed null curves of this family can be found in [8, Figures 2 and 3]).
Figure 1: Left: The closed null curve $\gamma$ with constant bending $\kappa_{7,3}$ (see Example 2.6). This curve represents a torus knot of type $(2,5)$. Right: The associated pair of star-shaped cousins $(\eta_{+},\eta_{-})$ in blue and red, respectively.
3 The $\boldsymbol{\mathcal{T}}$-transform
In this section, we will introduce a geometric transformation on null curves in ${\rm AdS}$, which we call the $\mathcal{T}$-transform, and show how to build it using a solution of the Riccati equation.
In order to define the $\mathcal{T}$-transform on null curves, we will employ Tabachnikov’s transformation [^2] In [10], a symmetric relation is introduced for circle maps by considering the continuous version of a relation for ideal polygons in the hyperbolic plane. Extending this relation to the lifts of these circle maps to $\mathbb{R}^{2}$, a transformation is defined. This transformation is, essentially, our $\mathcal{T}$-transform on star-shaped curves. on star-shaped curves [10] and translate it to null curves through the associated pair of cousins.